# TSINR: Capturing Temporal Continuity via Implicit Neural Representations for Time Series Anomaly Detection

### v2

@@ Appendix D Study on the Group Number
Group number is a hyper-parameter in our method, representing the number of groups in the group-based architecture. In other words, it determines the number of variables learned by neurons in each group layer. Table 4 demonstrates the ablation studies on the group number. When the group number is equal to 1, it means a regular function is used rather than a group-based architecture. We observe that different datasets have different optimal parameters. This is reasonable. Drawing an analogy to image data, partitioning can accelerate the fitting of INR for images (Liu et al., 2023a). Partitioning is similar to the group-based architecture discussed here. For each image, the number of partitions is not fixed and depends on the inherent characteristics of each signal. Therefore, in our approach, the optimal group number parameter for each dataset also depends on the intrinsic features of that dataset. Nevertheless, using the group-based architecture has enhanced the model’s ability for anomaly detection.
Appendix E Study on the Anomaly Proportion
Appendix F Comparison to Other Classical Anomaly Detection Methods
Besides reconstruction-based methods, we also compared our approach with other classic anomaly detection methods, such as clustering-based methods (DeepSVDD (Ruff et al., 2018) and ITAD (Shin et al., 2020)), the density-estimation models (LOF (Breunig et al., 2000), MPPCACD (Yairi et al., 2017), DAGMM (Zong et al., 2018)), the classic methods (OCSVM (Tax and Duin, 2004), iForest (Liu et al., 2008)), and the change point detection and time series segmentation methods (BOCPD (Adams and MacKay, 2007), U-Time (Perslev et al., 2019), TS-CP2 (Deldari et al., 2021)). As shown in Table 6, TSINR achieves comparable performance on SMAP dataset and outperforms other baselines on other datasets.
Appendix G Full Experiment Results on UCR Dataset
Due to the page limitation, we place the full experimental results conducted on UCR data in this section. The UCR dataset mainly involves four domains, including human medicine, meteorology, biology and industry (Wu and Keogh, 2021). We follow the settings of the previous work (Rewicki et al., 2023) to categorize the 250 sub-datasets into four groups, with each group corresponding to a specific domain. Figure 9 illustrates the average results within each category. It can be observed that our method achieves the best performance in each domain, demonstrating its strong generalizability.
Appendix H Efficiency Analysis
Figure 8. The kernel density estimation plots of five multivariate datasets.
@@ Appendix H Efficiency Analysis
Dataset SMD PSM MSL
Metric P R F1 P R F1 P R F1
$\gamma$ = 0.5 83.09 80.46 81.76 99.53 88.65 93.78 94.67 61.92 74.87
$\gamma$ = 0.6 80.55 80.93 80.74 99.37 88.90 93.85 93.16 62.36 74.71
$\gamma$ = 0.7 77.45 81.46 79.40 99.33 89.00 93.89 92.32 66.77 77.49
$\gamma$ = 0.8 75.16 83.06 78.91 99.27 88.80 93.75 91.92 73.43 81.64
$\gamma$ = 0.9 73.44 83.37 78.09 98.82 89.03 93.67 90.46 75.20 82.12
$\gamma$ = 1.0 72.55 83.39 77.59 99.21 89.37 94.04 83.98 84.26 84.12
Table 6. Comparison to other classical anomaly detection methods. The precision (P), recall (R), and F1-score (F1) values are reported, all in percentage (%). The best results are in Bold and the second-best results are in underlined.
Dataset SMD PSM SWaT MSL SMAP Average
Metric P R F1 P R F1 P R F1 P R F1 P R F1 F1
LOF 56.34 39.86 46.68 57.89 90.49 70.61 72.15 65.43 68.62 47.72 85.25 61.18 58.93 56.33 57.6 60.94
OCSVM 44.34 76.72 56.19 62.75 80.89 70.67 45.39 49.22 47.23 59.78 86.87 70.82 53.85 59.07 56.34 60.25
U-Time 65.95 74.75 70.07 82.85 79.34 81.06 46.2 87.94 60.58 57.2 71.66 63.62 49.71 56.18 52.75 65.62
IForest 42.31 73.29 53.64 76.09 92.45 83.48 49.29 44.95 47.02 53.94 86.54 66.45 52.39 59.07 55.53 61.22
DAGMM 67.3 49.89 57.3 93.49 70.03 80.08 89.92 57.84 70.4 89.6 63.93 74.62 86.45 56.73 68.51 70.18
ITAD 86.22 73.71 79.48 72.8 64.02 68.13 63.13 52.08 57.08 69.44 84.09 76.07 82.42 66.89 73.85 70.92
VAR 78.35 70.26 74.08 90.71 83.82 87.13 81.59 60.29 69.34 74.68 81.42 77.9 81.38 53.88 64.83 74.66
MMPCACD 71.2 79.28 75.02 76.26 78.35 77.29 82.52 68.29 74.73 81.42 61.31 69.95 88.61 75.84 81.73 75.74
CL-MPPCA 82.36 76.07 79.09 56.02 99.93 71.8 76.78 81.5 79.07 73.71 88.54 80.44 86.13 63.16 72.88 76.66
TS-CP2 87.42 66.25 75.38 82.67 78.16 80.35 81.23 74.1 77.5 86.45 68.48 76.42 87.65 83.18 85.36 79.00
Deep-SVDD 78.54 79.67 79.1 95.41 86.49 90.73 80.42 84.45 82.39 91.92 76.63 83.58 89.93 56.02 69.04 80.97
TSINR 83.09 80.46 81.76 99.21 89.37 94.04 99.31 72.32 83.69 83.57 85.40 84.47 91.67 76.42 83.35 85.46
Table 7. The efficiency of TSINR on the data with 128 batch size.
0.1s 0.08s 8.3M
Figure 9. We measure the average Precision (Left), Recall (Middle), and F1 score (Right) among different fields in the UCR datasets, which demonstrates that our TSINR is robust to time series data from different fields.

### v3

@@ Appendix D Study on the Group Number
The group number is a hyperparameter in our method, indicating the number of groups in the group-based architecture. It determines the number of variables learned by neurons in each group layer. Table 8 presents ablation studies on the group number. When set to 1, the method reverts to a regular function without grouping. Our results show that the optimal group number varies across datasets, which is reasonable given that, similar to image data, partitioning accelerates INR fitting and is dependent on the dataset’s intrinsic characteristics. Nonetheless, the group-based architecture improves the model’s anomaly detection performance.
Table 5. The additional results on multivariate and univariate datasets. The AUC and VUS values are reported. The best results are in Bold and the second best ones are underlined.
SMD AUC 0.747 0.654 0.765 0.652 0.728 0.729 0.760 0.738 0.732 0.766 0.723 0.774
VUS 0.740 0.639 0.762 0.631 0.723 0.722 0.755 0.734 0.728 0.762 0.720 0.769
PSM AUC 0.721 0.662 0.666 0.661 0.704 0.712 0.622 0.585 0.565 0.590 0.578 0.722
VUS 0.667 0.563 0.608 0.556 0.657 0.670 0.609 0.570 0.543 0.575 0.568 0.674
SWaT AUC 0.816 0.817 0.820 0.817 0.818 0.816 0.444 0.737 0.622 0.247 0.236 0.823
VUS 0.518 0.512 0.530 0.521 0.534 0.515 0.435 0.701 0.598 0.241 0.232 0.754
MSL AUC 0.624 0.550 0.532 0.550 0.602 0.613 0.596 0.601 0.615 0.623 0.590 0.657
VUS 0.607 0.525 0.518 0.525 0.569 0.599 0.555 0.569 0.580 0.591 0.552 0.638
SMAP AUC 0.526 0.450 0.456 0.450 0.452 0.490 0.401 0.380 0.397 0.455 0.474 0.576
VUS 0.502 0.418 0.450 0.415 0.438 0.482 0.363 0.343 0.373 0.412 0.444 0.567
PTB-XL AUC 0.604 0.485 0.583 0.485 0.536 0.560 0.589 0.545 0.516 0.618 0.627 0.660
VUS 0.456 0.339 0.436 0.339 0.386 0.417 0.453 0.401 0.365 0.471 0.486 0.534
SKAB AUC 0.536 0.429 0.493 0.440 0.570 0.496 0.482 0.480 0.504 0.496 0.496 0.585
VUS 0.535 0.421 0.492 0.434 0.569 0.495 0.482 0.474 0.504 0.496 0.496 0.585
UCR AUC 0.520 0.541 0.548 0.533 0.546 0.517 0.502 0.523 0.546 0.533 0.529 0.663
VUS 0.511 0.513 0.538 0.508 0.533 0.506 0.486 0.511 0.528 0.514 0.513 0.650
Dataset SMD PSM MSL
Metric P R F1 P R F1 P R F1
$\gamma$ = 0.5 83.09 80.46 81.76 99.53 88.65 93.78 94.67 61.92 74.87
$\gamma$ = 0.6 80.55 80.93 80.74 99.37 88.90 93.85 93.16 62.36 74.71
$\gamma$ = 0.7 77.45 81.46 79.40 99.33 89.00 93.89 92.32 66.77 77.49
$\gamma$ = 0.8 75.16 83.06 78.91 99.27 88.80 93.75 91.92 73.43 81.64
$\gamma$ = 0.9 73.44 83.37 78.09 98.82 89.03 93.67 90.46 75.20 82.12
$\gamma$ = 1.0 72.55 83.39 77.59 99.21 89.37 94.04 83.98 84.26 84.12
Table 7. The efficiency of TSINR on the data with 128 batch size.
0.1s 0.08s 8.3M